A complex number z satisfies the equation z^2=3 +4i which of the following is true about z?


1
Expert's answer
2021-10-25T19:22:11-0400

Complete question




"(3+4i)^2 = (3+4i)(3+4i) =\\
9 + 12i + 12i + 16i² = \\
9 + 24i+16(-1) =\\
9 -16 +24i=\\
-7i+24i"


"\therefore" The correct option is the last option"."